OpenStudy (anonymous):

Medal + Fan Your friend’s job pays $10 per hour and he works 40 hours per week. His tax rate is 20%. Assuming he works 50 weeks of the year his yearly gross income is A $4,000 B $16,000 C $20,000 D $40,000

OpenStudy (anonymous):

@Abhisar @arabpride @Auxuris @DangerousJesse @DangerousJesse @Eklipsed @HeatDude @Haseeb96 @hotdancingbabe14 @gorv @kl0723 @Librarian @linn99123 @CrazyTurtle483 @undeadknight26 @Phebe

OpenStudy (anonymous):

10 x 40 = 400 x .20(tax) = 80 400 - 80 = 320 x 50 = 16,000 B is your answer

OpenStudy (librarian):

I told you guys and girls I don't really specialize in math /=

OpenStudy (anonymous):

can you hel me with this question?@HeatDude

OpenStudy (anonymous):

@Basher @zepdrix @satellite73

OpenStudy (gorv):

payment=10/hour this ia payment of one hour in oneweek he works 40 hour so 1 week pay=40*10=400/week in a year he wprks for 50 week

OpenStudy (gorv):

so 50 week pay= 400*50=20000 now tax = 20% of total income tax=20000*20/100=4000 so subtract tax value for gross income 20000-4000=16000